OpenStudy (anonymous):

Evaluate log_4(1/cuberoot(16))

OpenStudy (amoodarya):

\[\log_{4} \sqrt[3]{16}= ?\]

OpenStudy (anonymous):

I feel I have to use log_bB^x=x, but I just can't get it. The answer is supposed to be -2/3. But as far as I've gotten is: log_4(16^-1/3) and i have the feeling I should put 16 as 4^2.. but something just doesn't work out

OpenStudy (amoodarya):

\[\log_{b^n}a^m=\frac{ m }{ n } \log_{b} a\]

OpenStudy (anonymous):

@amoodarya yes exactly that

OpenStudy (anonymous):

@amoodarya How would I apply it? I see the base is raised to the n power, do i make it a 2^2?

OpenStudy (amoodarya):

\[\sqrt[3]{16}=2^{\frac{ 4 }{ 3 }}\]

OpenStudy (amoodarya):

\[\log_{4} \frac{1}{\sqrt[3]{16}}=x\\4^x=\frac{1}{\sqrt[3]{16}}\\2^{2x}=2^{\frac{ -4 }{ 3 }}\\2x=\frac{ -4 }{ 3 }\]

OpenStudy (anonymous):

Thanks for the help!